# Rigidity of proper quasi-homogeneous domains in positive flag manifolds

### v1

@@ 2.2.2. Cartan decomposition
For any $\alpha\in\Sigma$ and $X\in\mathfrak{g}_{\alpha}\smallsetminus\{0\}$, there exists a unique scalar multiple $X^{\prime}$ of $X$ such that $(X^{\prime},\sigma(-X^{\prime}),[\sigma(X^{\prime}),X^{\prime}])$ is an $\mathfrak{sl}_{2}$-triple. The element $[\sigma(X^{\prime}),X^{\prime}]$ does not depend on the choice of $X\in\mathfrak{g}_{\alpha}$, and is denoted by $h_{\alpha}$. The family $(h_{\alpha})_{\alpha\in\operatorname{\Delta}}$, forms a basis of $\mathfrak{a}$, whose dual basis in $\mathfrak{a}^{*}$ is denoted by $(\omega_{\alpha})_{\alpha\in\operatorname{\Delta}}$.
@@ 2.2.4. The restricted Weyl group
(2.2) $\operatorname{Lie}(P_{\Theta})=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma^{+}}\oplus\bigoplus_{\alpha\in\Sigma\smallsetminus\Sigma_{\Theta}^{+}}\mathfrak{g}_{-\alpha}.$
@@ 2.3. Flag manifolds
Let $G$ be a real semisimple Lie group. The flag manifold associated with $\Theta$ is the quotient space $\mathscr{F}_{\Theta}:=G/P_{\Theta}$. The flag manifold opposite to $\mathscr{F}_{\Theta}$ is $\mathscr{F}_{\Theta}^{-}:=G/P_{\Theta}^{-}$. The action of $G$ on $\mathscr{F}_{\Theta}\times\mathscr{F}_{\Theta}^{-}$ by left translations has exactly one open orbit $\mathscr{O}$, which is the orbit of $(P_{\Theta},P^{-}_{\Theta})$ and is dense. Two elements $x\in\mathscr{F}_{\Theta}$ and $y\in\mathscr{F}_{\Theta}^{-}$ are said to be transverse if $(x,y)\in\mathscr{O}$.
@@ 3.3. An invariant cone and causality
Let $x\in\mathbb{A}$ and $g\in G$ be such that $g\cdot P\in\mathbb{A}$. Then for any $\delta_{1}\in\{-,+\}$, there exist $\delta_{2}\in\{-,+\}$ and a neighborhood $\mathscr{U}$ of $x$ such that $g\cdot(\mathscr{U}\cap\mathbf{I}^{\delta_{1}}(x))\subset\mathbf{I}^{\delta_{2}}(g\cdot x)$.
@@ 4. Proper domains and their automorphism group
In this section we recall some definitions and properties of domains in a flag manifold. In particular, we introduce the diamonds (see Definition 4.6) in $\mathbf{Sb}(G)$, and recall their basic properties. These proper domains are quasi-homogeneous, and Theorem 1.4 states that they are the only proper domains in $\mathbf{Sb}(G)$ with this property.
@@ 4.1. Generalities on proper domains
A domain $\Omega$ is said to be quasi-homogeneous if there exists a compact subsets $\mathscr{K}\subset\Omega$ such that $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$. Note that if $G$ is an $\operatorname{HTT}$ Lie group, then this notion only depends on the Lie algebra $\mathfrak{g}$ of $G$, because $\operatorname{Ad}(G)$ has finite index in $\mathsf{Aut}(\mathfrak{g})$ (see e.g. [Sat14]). The domain $\Omega$ is said to be symmetric if for any $x\in\Omega$ there exists an order-two element $s_{x}\in\mathsf{Aut}_{1}(\mathfrak{g})$ such that $s_{x}\cdot\Omega=\Omega$ and $x$ is the only fixed point of $s_{x}$ in $\Omega$. This notion only depends on the Lie algebra $\mathfrak{g}$ of $G$ (recall that, if $G$ is an $\operatorname{HTT}$ Lie group, then one has $\mathsf{Aut}_{1}(\mathfrak{g})=\mathsf{Aut}(\mathfrak{g})$ — see Section 3.1).
Note that $\mathsf{Aut}(g\cdot\Omega)=g\mathsf{Aut}(\Omega)g^{-1}$ for all $g\in G$; therefore the property of being quasi-homogeneous, resp. symmetric, is invariant under the action of $G$ on $\mathscr{F}_{\Theta}$. In our case, i.e. when $G$ is an $\operatorname{HTT}$ Lie group, by Remark 4.2 and Section 3.2, it will always be possible to assume that $\Omega$ is proper in $\mathbb{A}$, and given a point $x\in\overline{\Omega}$, we can always assume that $x=P$, up to translating $\Omega$ by an element of $G$.
@@
Let $\Omega\subset\mathscr{F}_{\Theta}$ be a domain. The dual of $\Omega$ is the set
@@
The domain $\Omega$ is dually convex, if for any $p\in\partial\Omega$ there exists $\xi\in\Omega^{*}$ such that $p\in\operatorname{Z}_{\xi}$. One has:
Proposition 4.4.
([Zim18a, Cor. 9.3]) Any proper quasi-homogeneous domain of $\mathscr{F}_{\Theta}$ is dually convex.
@@ 4.2. Diamonds
    Diamonds are defined more generally in any flag manifold $\mathscr{F}_{\Theta}$ admitting a $\Theta$-positive structure [GLW21]. By Zimmer’s theorem (Fact 1.3) and Fact 4.8, these diamonds are quasi-homogeneous (resp. divisible) if, and only if, the flag manifold $\mathscr{F}_{\Theta}$ is the Shilov boundary associated with an $\operatorname{HTT}$ Lie group.
@@ 5.1. Definition and basic properties
The following lemma follows from Lemma 3.8:
@@
Remark 5.3.
Recall that a Schubert subvariety of $\mathbf{Sb}(G)$ is the Zariski closure in $\mathbf{Sb}(G)$ of a subset of the form $X(w)=P_{\operatorname{\Delta}}w\cdot P$, where $\omega$ is an element of the Weyl group of $\Sigma$ stabilizing $\alpha_{r}$ (see e.g. [Kna96]). Noticing that $\Lambda_{\mathsf{std}}$ is a one-dimensional Schubert subvariety of $\mathbf{Sb}(G)$ gives a direct proof of Lemma 5.2.
@@ 6. An invariant metric
In this section we define the Kobayashi metric of a proper domain $\Omega$ of $\mathbf{Sb}(G)$. Constructions of Kobayashi metrics are classical and were initiated by S. Kobayashi [Kob67, Sho84]. The properties of the Kobayashi metric (in particular, its properness, see Corollary 6.13) in the quasi-homogeneous case will allow us to relate the geometry of the boundary of a proper quasi-homogeneous domain to the dynamics of its automorphism group in Section 7.
@@ 6.1. A Kobayashi metric
Let $x,y\in\Omega$ be two conjugate points. Let $\Lambda_{x,y}\in\mathscr{L}$ be the unique photon containing $x$ and $y$, and let $g\in G$ be such that $\Lambda_{x,y}=g\cdot\Lambda_{\mathsf{std}}$. We denote by $I_{x,y}$ the connected component of $\Omega\cap\Lambda_{x,y}$ containing $x$ and $y$. Recall the parametrization $\mathsf{i}_{g}:\mathbb{P}(\mathbb{R}^{2})\rightarrow\Lambda_{x,y}$ of (5.1). We define:
@@ 6.2.1. The Caratheodory metric
After this paper was completed, we learnt that Beyrer–Guichard–Labourie–Pozzetti–Wienhard [BGL+] prove a similar property for flag manifolds with a $\Theta$-positive structure. Other versions are established in [Zim15, Lem. 10.4] for Grassmannians $\operatorname{Gr}_{p}(\mathbb{R}^{n})$ and in [LZ17, Lem. 2.9] for the full flag manifold of $\operatorname{SL}(n,\mathbb{R})$.
@@
6.2.3. Properness of $K_{\Omega}$.
@@
If $\Omega\subset\mathbf{Sb}(G)$ is a proper dually convex domain, then $K_{\Omega}$ is a proper metric. In particular, if $\Omega\subset\mathbf{Sb}(G)$ is a proper quasi-homogeneous domain, then $K_{\Omega}$ is a proper metric.
Proposition 6.8 and Corollary 6.13 imply in particular that $K_{\Omega}$ is geodesic as soon as $\Omega$ is a proper quasi-homogeneous (and even just dually convex) domain, although we will not use this fact.
@@
The only consequence of Proposition 6.8 and Corollary 6.13 we will use in the proof of Theorem 1.4 is the fact that $K_{\Omega}$ is proper as soon as $\Omega$ is quasi-homogeneous, and that the $K_{\Omega}$-length of a $1$-chain $(x,y)$ is in this case equal to $\mathsf{k}_{\Omega}(x,y)$. Other approaches would have led to this conclusion: for instance, the theory of Euclidean Jordan algebras, or looking at explicit proximal representations of $G$ for each value of $\mathfrak{g}$ in Table 1. However, the approach we choose makes it possible to compare the Kobayashi metric with the already existing Caratheodory metrics, and is generalizable to other flag manifolds in which Kobayashi constructions have already been made (see Remark 1.10).
@@ 7. The dynamics at the boundary
This section is divided into two parts. Section 7.1 is devoted to the definition and the analysis of two families of points of the boundary of a proper domain in $\mathbf{Sb}(G)$, those of visual and $\mathscr{R}$-extremal points of the domain. In Section 7.2, we prove the existence of $\mathscr{R}$-extremal points satisfying a particular geometric property, namely strongly $\mathscr{R}$-extremal points.
7.1. Visual points and $\mathscr{R}$-extremal points
In this section, following [LZ19], we investigate the relation between the structure of the boundary of a proper quasi-homogeneous domain and the dynamics of its automorphism group.
7.1.1. Visual points
Let $\Omega\subset\mathbf{Sb}(G)$ be a proper domain. A point $p\in\partial\Omega$ will be called a visual point if it satisfies the following condition: for any pair of sequences $(x_{k}),(y_{k})$ of $\Omega$ such that $(x_{k})$ converges to $p$, if the sequence $(K_{\Omega}(x_{k},y_{k}))$ is bounded in $\mathbb{R}_{\geq 0}$, then $y_{k}\rightarrow p$.
Visual points satisfy the following geometric property:
Lemma 7.1.
Assume that $\Omega$ is a proper quasi-homogeneous domain of $\mathbf{Sb}(G)$. Let $p\in\partial\Omega$ be a visual point. Then $\operatorname{Z}_{p}\cap\Omega=\emptyset$.
  • Proof.
    Since $\Omega$ is quasi-homogeneous we can find $x\in\Omega$ and a sequence $(x_{k})\in\Omega^{\mathbf{N}}$ converging to $x$ such that, for some sequence $(g_{k})\in\mathsf{Aut}(\Omega)^{\mathbb{N}}$,
    $g_{k}\cdot x_{k}\underset{k\rightarrow+\infty}{\longrightarrow}p.$
    Since $p$ is visual, one has $g_{k}\cdot x\rightarrow p$. By invariance of $K_{\Omega}$ and the fact that $p$ is visual, one actually has $g_{k}\cdot y\rightarrow p$ for all $y\in\Omega$.
    Let $B\subset\Omega$ be the closed ball of center $x_{0}\in\Omega$ and of radius $1$ for the metric $K_{\Omega}$. If the sequence $g_{k}\cdot B$ does not converge to $\{p\}$ for the Hausdorff topology, then there is a neighborhood $\mathscr{V}$ of $p$ in $\mathbf{Sb}(G)$ and a sequence $(y_{k})\in B^{\mathbb{N}}$ such that $g_{k}\cdot y_{k}\notin\mathscr{V}$ for all $k\in\mathbb{N}$. Since $B$ is a compact subset of $\Omega$ by Corollary 6.13, up to extracting we may assume that there exists $y\in B$ such that $g_{k}\cdot y_{k}\rightarrow y$. Then $(g_{k}\cdot y)$ converges to $p$. But $K_{\Omega}(g_{k}\cdot y_{k},g_{k}\cdot y)\rightarrow 0$, and in particular the sequence $\left(K_{\Omega}(g_{k}\cdot y_{k},g_{k}\cdot y)\right)$ is bounded, so by visuality $g_{k}\cdot y_{k}\rightarrow p$, contradiction.
    Hence $g_{k}\cdot B\rightarrow\{x\}$ for the Hausdorff topology. Then, by the KAK decomposition, up to extracting, we may assume that there exists $q\in\mathbf{Sb}(G)$ such that $(g_{k})$ is $(p,q)$-contracting, in the sense that
    $g_{k|\mathbf{Sb}(G)\smallsetminus\operatorname{Z}_{q}}\rightarrow p,$
[…]
Remark 7.2.
(1) Lemma 7.1 is specific to our context: given a point $p$ of an arbitrary flag manifold $\mathscr{F}_{\Theta}$, the set $\operatorname{Z}_{p}$ is a subset of $\mathscr{F}_{\Theta}^{-}$, which cannot be $G$-equivariantly identified with $\mathscr{F}_{\Theta}$ in general. There are versions of Lemma 7.1 for flag manifolds that are not self-opposite, but they express a weaker geometric property for visual points (see e.g. [Gal25]). For instance, for the flag manifold $\mathbb{P}(\mathbb{R}^{n})$ with $n\geq 2$, which is as far as possible from being self-opposite, this weaker version of Lemma 7.1 expresses that any visual point of $\partial\Omega$ is contained in a supporting projective hyperplane to $\Omega$, which is already a consequence of the convexity of $\Omega$.
(2) The name “visual” refers to the visual boundary of the abstract metric space $(\Omega,K_{\Omega})$, defined as the set of all unbounded sequences of $(\Omega,K_{\Omega})$, quotiented by the relation
$(x_{k})\ \mathscr{R}\ (y_{k})\Longleftrightarrow(K_{\Omega}(x_{k},y_{k}))\text{ is bounded}.$
There is a natural embedding from the set of visual points of $\partial\Omega$ to the visual boundary of $(\Omega,K_{\Omega})$.
7.1.2. $\mathscr{R}$-extremal points

### v2

@@ 2.2.2. Cartan decomposition
For any $\alpha\in\Sigma$ and $X\in\mathfrak{g}_{\alpha}\smallsetminus\{0\}$, there exists a unique scalar multiple $X^{\prime}$ of $X$ such that $(X^{\prime},\sigma(X^{\prime}),X^{\prime}],\sigma(-X^{\prime}))$ is an $\mathfrak{sl}_{2}$-triple. The element $[\sigma(X^{\prime}),X^{\prime}]$ does not depend on the choice of $X\in\mathfrak{g}_{\alpha}$, and is denoted by $h_{\alpha}$. The family $(h_{\alpha})_{\alpha\in\operatorname{\Delta}}$, forms a basis of $\mathfrak{a}$, whose dual basis in $\mathfrak{a}^{*}$ is denoted by $(\omega_{\alpha})_{\alpha\in\operatorname{\Delta}}$.
@@ 2.2.4. The restricted Weyl group
(2.2) $\operatorname{Lie}(P_{\Theta})=\mathfrak{g}_{0}\oplus\bigoplus_{\alpha\in\Sigma^{+}}\mathfrak{g}_{\alpha}\oplus\bigoplus_{\alpha\in\Sigma\smallsetminus\Sigma_{\Theta}^{+}}\mathfrak{g}_{-\alpha}.$
@@ 2.3. Flag manifolds
Let $G$ be a real semisimple Lie group. The flag manifold associated with $\Theta$ is the quotient space $\mathscr{F}_{\Theta}:=G/P_{\Theta}$. The flag manifold opposite to $\mathscr{F}_{\Theta}$ is $\mathscr{F}_{\Theta}^{-}:=G/P_{\Theta}^{-}$.
2.3.1. Transversality
The action of $G$ on $\mathscr{F}_{\Theta}\times\mathscr{F}_{\Theta}^{-}$ by left translations has exactly one open orbit $\mathscr{O}$, which is the orbit of $(P_{\Theta},P^{-}_{\Theta})$ and is dense. Two elements $x\in\mathscr{F}_{\Theta}$ and $y\in\mathscr{F}_{\Theta}^{-}$ are said to be transverse if $(x,y)\in\mathscr{O}$.
@@ 3.3. An invariant cone and causality
Let $x\in\mathbb{A}$ and $g\in G$ be such that $g\cdot x\in\mathbb{A}$. Then for any $\delta_{1}\in\{-,+\}$, there exist $\delta_{2}\in\{-,+\}$ and a neighborhood $\mathscr{U}$ of $x$ such that $g\cdot(\mathscr{U}\cap\mathbf{I}^{\delta_{1}}(x))\subset\mathbf{I}^{\delta_{2}}(g\cdot x)$.
@@ 4. Proper domains and their automorphism group
In this section we recall some definitions and properties of domains in a flag manifold. In particular, we introduce the diamonds (see Definition 4.6) in $\mathbf{Sb}(G)$, and recall their basic properties. These proper domains are almost-homogeneous, and Theorem 1.4 states that they are the only proper domains in $\mathbf{Sb}(G)$ with this property.
@@ 4.1. Generalities on proper domains
The domain $\Omega$ is said to be almost-homogeneous is there exists a compact subset $\mathscr{K}\subset\Omega$ such that $\Omega=\mathsf{Aut}(\Omega)\cdot\mathscr{K}$.
The full orbital limit set of $\Omega$ is the set
$\Lambda_{\Omega}^{\operatorname{orb}}:=\bigcup_{x\in\Omega}(\overline{\mathsf{Aut}(\Omega)\cdot x})\smallsetminus(\mathsf{Aut}(\Omega)\cdot x),$
see [DGK24]. Since $\mathsf{Aut}(\Omega)$ acts properly on $\Omega$, we have $\Lambda_{\Omega}^{\operatorname{orb}}\subset\partial\Omega$. A domain $\Omega$ is said to be almost-homogeneous if $\Lambda_{\Omega}^{\operatorname{orb}}=\partial\Omega$. This is equivalent to saying that for all $p\in\partial\Omega$, there exist $x\in\Omega$ and $(g_{n})\in\mathsf{Aut}(\Omega)^{\mathbb{N}}$ such that $g_{n}\cdot x\rightarrow p$, and for proper domains it is a priori weaker than quasi-homogeneity. Note that if $G$ is an $\operatorname{HTT}$ Lie group, then the notions of quasi-homogeneity and almost-homogeneity only depends on the Lie algebra $\mathfrak{g}$ of $G$, because $\operatorname{Ad}(G)$ has finite index in $\mathsf{Aut}(\mathfrak{g})$ (see e.g. [Sat14]). The domain $\Omega$ is said to be symmetric if for any $x\in\Omega$ there exists an order-two element $s_{x}\in\mathsf{Aut}_{1}(\mathfrak{g})$ such that $s_{x}\cdot\Omega=\Omega$ and $x$ is the only fixed point of $s_{x}$ in $\Omega$. This notion only depends on the Lie algebra $\mathfrak{g}$ of $G$ (recall that, if $G$ is an $\operatorname{HTT}$ Lie group, then one has $\mathsf{Aut}_{1}(\mathfrak{g})=\mathsf{Aut}(\mathfrak{g})$ — see Section 3.1).
Note that $\mathsf{Aut}(g\cdot\Omega)=g\mathsf{Aut}(\Omega)g^{-1}$ for all $g\in G$; therefore the property of being almost-homogeneous, resp. symmetric, is invariant under the action of $G$ on $\mathscr{F}_{\Theta}$. In our case, i.e. when $G$ is an $\operatorname{HTT}$ Lie group, by Remark 4.2 and Section 3.2, it will always be possible to assume that $\Omega$ is proper in $\mathbb{A}$, and given a point $x\in\overline{\Omega}$, we can always assume that $x=P$, up to translating $\Omega$ by an element of $G$.
@@
We will make use of the following lemma:
Lemma 4.4.
Let $\Omega,\Omega^{\prime}$ be two proper domains of $\mathscr{F}_{\Theta}$ such that $\Omega\subset\Omega^{\prime}$. Assume that $\mathsf{Aut}(\Omega)\subset\mathsf{Aut}(\Omega^{\prime})$ and that $\Omega$ is almost-homogeneous. Then $\Omega=\Omega^{\prime}$.
  • Proof.
Let $p\in\partial\Omega$. There exist $x\in\Omega$ and $(g_{n})\in\mathsf{Aut}(\Omega)^{\mathbb{N}}$ such that $g_{n}\cdot x\rightarrow p$. Thus $(g_{n})$ is unbounded in $G$. Since $\mathsf{Aut}(\Omega)\subset\mathsf{Aut}(\Omega^{\prime})$, the group $\mathsf{Aut}(\Omega)$ acts properly on the proper domain $\Omega^{\prime}$. Thus $p\in\partial\Omega^{\prime}$.
[…]
We can now prove Lemma 1.2. The proof is similar to that of [Zim18a]. We give it for convenience:
  • Proof of Lemma 1.2.
    Let us denote by $\Omega_{i}\subset\mathbf{Sb}(G)$ the image of $\Omega$ by the canonical projection $\Pi_{i}:\mathbf{Sb}(G)\rightarrow\mathbf{Sb}(G_{i})$ for $1\leq i\leq k$, and let $\Omega^{\prime}:=\Omega_{1}\times\cdots\times\Omega_{k}$. Then $\Omega^{\prime}$ is a proper $\mathsf{Aut}(\Omega)$-invariant domain of $\mathbf{Sb}(G)$ containing $\Omega$. Then by Lemma 4.4, we have $\Omega=\Omega^{\prime}$.
[…]
@@
Let $\Omega\subset\mathscr{F}_{\Theta}$ be a subset. The dual of $\Omega$ is the set
@@
The domain $\Omega$ is dually convex, if for any $p\in\partial\Omega$ there exists $\xi\in\Omega^{*}$ such that $p\in\operatorname{Z}_{\xi}$.
In [Zim18a, Cor. 9.3], A. Zimmer proves that any proper almost-homogeneous domain of $\mathscr{F}_{\Theta}$ is dually convex. Based on his proof, we give in Section 6.2.1 a slightly stronger result (see Proposition 6.9).
@@
  2. (2)
    If $\Omega\subset\mathscr{F}_{\Theta}$ is a proper domain, then its bidual $\Omega^{**}$ is always a proper dually convex domain of $\mathscr{F}_{\Theta}$ by definition.
@@ 4.2. Diamonds
    Diamonds are defined more generally in any flag manifold $\mathscr{F}_{\Theta}$ admitting a $\Theta$-positive structure [GLW21]. By Zimmer’s theorem (Fact 1.3) and Fact 4.8, these diamonds are almost-homogeneous (resp. divisible) if, and only if, the flag manifold $\mathscr{F}_{\Theta}$ is the Shilov boundary associated with an $\operatorname{HTT}$ Lie group.
@@ 5.1. Definition and basic properties
The following lemma is a direct consequence of Lemma 3.8:
@@ 6. An invariant metric
In this section we define the Kobayashi metric of a proper domain $\Omega$ of $\mathbf{Sb}(G)$. Constructions of Kobayashi metrics are classical and were initiated by S. Kobayashi [Kob67, Sho84]. The properties of the Kobayashi metric (in particular, its properness, see Corollary 6.15) in the almost-homogeneous case will allow us to relate the geometry of the boundary of a proper almost-homogeneous domain to the dynamics of its automorphism group in Section 7.
@@ 6.1. A Kobayashi metric
Let $x,y\in\Omega$ be two conjugate points. Let $\Lambda_{x,y}\in\mathscr{L}$ be a photon containing $x$ and $y$, and let $g\in G$ be such that $\Lambda_{x,y}=g\cdot\Lambda_{\mathsf{std}}$. We denote by $I_{x,y}$ the connected component of $\Omega\cap\Lambda_{x,y}$ containing $x$ and $y$. Recall the parametrization $\mathsf{i}_{g}:\mathbb{P}(\mathbb{R}^{2})\rightarrow\Lambda_{x,y}$ of (5.1). We define:
@@ 6.1.3. Kobayashi hyperbolicity
Remark 6.8.
The proof of Proposition 6.7 gives that for any two sequences $(x_{k}),(y_{k})\in\Omega^{\mathbb{N}}$ such that $x_{k}\rightarrow p\in\partial\overline{\Omega}$ and $y_{k}\rightarrow q\in\overline{\Omega}$, if $K_{\Omega}(x_{k},y_{k})\rightarrow 0$ then one has $p=q$.
@@ 6.2.1. The Caratheodory metric
In [Zim18a], this metric is used to prove that quasi-homogeneous domains are dually convex. In this paper, we need the slightly stronger following result, whose proof relies on the one of [Zim18a, Cor. 9.3]:
Proposition 6.9.
Any proper almost-homogeneous domain of $\mathscr{F}_{\Theta}$ is dually convex.
  • Proof.
[…]
@@
After this paper was completed, we learnt that Beyrer–Guichard–Labourie–Pozzetti–Wienhard [BGL+24] prove a similar property for flag manifolds with a $\Theta$-positive structure. Other versions are established in [Zim15, Lem. 10.4] for Grassmannians $\operatorname{Gr}_{p}(\mathbb{R}^{n})$ and in [LZ17, Lem. 2.9] for the full flag manifold of $\operatorname{SL}(n,\mathbb{R})$.
@@
6.3. Properness
In this section we state a corollary of Proposition 6.10, whixh is the properness of the Kobayashi metric on a proper dually convex domain of $\mathbf{Sb}(G)$. This fact is of independant interest and will not be used in the rest of the paper.
@@ 6.3. Properness
If $\Omega\subset\mathbf{Sb}(G)$ is a proper dually convex domain, then $K_{\Omega}$ is a proper metric. In particular, if $\Omega\subset\mathbf{Sb}(G)$ is a proper almost-homogeneous domain, then $K_{\Omega}$ is a proper metric.
Proposition 6.10 and Corollary 6.15 imply in particular that $K_{\Omega}$ is geodesic as soon as $\Omega$ is a proper almost-homogeneous (and even just dually convex) domain, although we will not use this fact.
@@
The only consequence of Proposition 6.10 and Corollary 6.15 we will use in the proof of Theorem 1.4 is the fact that $K_{\Omega}$ is proper as soon as $\Omega$ is almost-homogeneous, and that the $K_{\Omega}$-length of a $1$-chain $(x,y)$ is in this case equal to $\mathsf{k}_{\Omega}(x,y)$. Other approaches would have led to this conclusion: for instance, the theory of Euclidean Jordan algebras, or looking at explicit proximal representations of $G$ for each value of $\mathfrak{g}$ in Table 1. However, the approach we choose makes it possible to compare the Kobayashi metric with the already existing Caratheodory metrics, and is generalizable to other flag manifolds in which Kobayashi constructions have already been made (see Remark 1.10).
@@ 7. The dynamics at the boundary
This section is divided into two parts. Section 7.1 is devoted to the definition and the analysis of a family of points of the boundary of a proper domain in $\mathbf{Sb}(G)$, those of $\mathscr{R}$-extremal points. In Section 7.2, we prove the existence of $\mathscr{R}$-extremal points satisfying a particular geometric property, namely strongly $\mathscr{R}$-extremal points.
7.1. $\mathscr{R}$-extremal points
In this section, following [LZ19], we investigate the relation between the structure of the boundary of a proper almost-homogeneous domain and the dynamics of its automorphism group.